Thermal capacity of 40 g of aluminium (s = 0.2 cal /g K) is       


        (a)    168 joule/°C      


(b)    672 joule/°C


(c)    840 joule/°C


(d)    33.6 joule/°C


Grade:11

2 Answers

Amit Saxena
10 years ago

(d)

Thermal capacity = ms = 40 × 0.2 = 8 Cal/°C

= 4.2 × 8 J = 33.6 joules/°C

 

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Thermal capacity = ms = 40 × 0.2
= 8 Cal/°C
= 4.2 × 8 J
= 33.6 joules/°C

Thanks and Regards